# SCREP: Scene Coordinate Regression and Evidential Learning-based Perception-Aware Trajectory Generation

### v1

@@ V-D Hardware-in-the-loop Experiment
Fig. 7(a) illustrates our hardware-in-the-loop experiment setup. The quadrotor flies inside an arena while real IMU data are logged and the onboard camera streams images to the virtual stata1 scene. The vehicle follows a rose-petal trajectory, $r=3\cos(2\theta)$, for three laps. Throughout the experiment, we compared the localization performance of the Forward baseline with that of our perception-aware trajectory planning framework.
In the experiment, our perception-aware trajectory planner performs better than the Forward baseline. Translation RMSE / mean error of the fused pose dropped from 26.3cm / 22.3cm to 20.7cm / 18.4cm, while rotational RMSE / mean error fell from 5.85 / 5.50 to 2.26 / 2.04. Fig. 7(b) shows the errors over time for one lap of tracking the rose petal trajectory. The time histories in Fig. 7(b) show that the error of our approach remains consistently below that of the baseline throughout a representative lap, confirming the effectiveness and feasibility of the proposed method in a real flight.
@@ VI Conclusion
This paper introduced a receding horizon, and perception‑aware trajectory generation framework that contains evidential learning-based scene coordinate regression (SCR). The optimizer consistently oriented the camera toward scene coordinates with low uncertainty and produced drift‑free, high‑accuracy poses in real-time. Simulation and hardware-in-the-loop experiments demonstrate that our framework significantly outperforms baseline methods in localization accuracy and is suitable for real-world flight. Notably, these gains were achieved even though the SCR network was trained with only a small numbers of images, underscoring the framework’s suitability for data‑scarce deployments. Future work will integrate obstacle avoidance, and test robustness to appearance changes such as lighting variations and object rearrangements.

### v2

@@ V-D Hardware-in-the-loop Experiment
The quadrotor follows a rose-petal trajectory, $r=3\cos(2\theta)$ for three laps during the hardware-in-the-loop experiment. Throughout the experiment, we compare the localization performance of the Forward baseline with our method.
Translation RMSE and mean error of the fused pose dropped from 29.7cm and 24.6cm to 22.7cm and 20.0cm, respectively, while rotational RMSE and mean error fell from 6.0 and 5.6 to 2.3 and 2.0. Fig. 7(b) shows the errors over time, demonstrating that the error of our approach remains consistently below that of the Forward throughout a representative lap of the rose petal trajectory. The results confirm the effectiveness and feasibility of the proposed method under close-to-real-world conditions.
@@ VI Conclusion
This paper introduces a receding-horizon, perception‑aware trajectory generation framework using evidential learning-based scene coordinate regression (SCR). Our algorithm consistently orients the camera toward scene coordinates with low uncertainty and produces drift‑free, high‑accuracy poses in real-time. Experiments demonstrate that our framework outperforms baseline methods in localization accuracy and is suitable for hardware-in-the-loop flight tests. Future work will integrate obstacle avoidance and test robustness to appearance changes such as lighting variations and object rearrangements.
ACKNOWLEDGMENTS
We used OpenAI ChatGPT for code design and debugging related to the experiments section, and for polishing and improving the clarity of writing across the entire paper. The authors reviewed, verified, and take full responsibility for all content.